Manuel is making a kite. he cuts out a piece of cloth in the shape of a diamond He uses two sticks as supports. The vertical stick is 90 cm long, and the horizontal stick is 80 cm long. They intersect at a right angle 60 cm from the bottom of the vertical stick. The vertical stick bisects the horizontal stick. Manuel wants to add a border around the perimeter of the kite. What is the length of material Manuel with need for the border? Show work and round to the nearest centimeter.
Answer by addingup(3677)

The vertical stick is 90. The two sticks meet at 60, so 90-60= 30. This means that you have two halves that are 30 long and two that are 60 long. This means four triangles (if you didn't draw it you should. I did and I'm looking at it). Each triangle has one side 40 and the others, as we just said, 30 and 60. The outside perimeter is formed by the lengths of the hypotenuses of these triangles. Find the hypotenuses, add them up, and you'll know the length of the border.
Top triangles:
30^2+40^2= 900+1600= 2500; sqrt2500= 50 and since we have two: 50*2= 100
Bottom:
60^2+40^2= 3600+1600= 5200; sqrt5200= 72; 72*2= 144
Perimeter:
100+144= 244 is the length of material